Mr.Ali has an income of 30000 and he purchase 100units. His income decreases by 20% and now he can purchase 90uni hits calculate income elasticity
1
Expert's answer
2020-04-29T09:11:29-0400

"Ei= \frac{ \Delta Q}{\Delta Income}" .


"\Delta Q = \frac{90-100}{100}=-10"%.

"\Delta Income = 20" %.

So, "Ei = \frac{-10}{20}=-0.5."